You are a shareholder in a C corporation. The corporation earns $ 2.01 per share before taxes. Once it has paid taxes it will distribute the rest of its earnings to you as a dividend. Assume the corporate tax rate is 40 % and the personal tax rate on​ (both dividend and​ non-dividend) income is 28 %. How much is left for you after all taxes are​ paid?
Answer #1

Profit before tax = $2.01 per share

Net Income per share = (1 - 0.40)2.01 = $1.206

Dividend = $1.206

After dividends are received,

After-tax amount = 1.206(1 - 0.28)

After-tax amount = $0.87